3x + 4y = 5
2x + 3y = 4
Solve this problem using elimination.
I wasn't sure how to do it, because you can't elimate any one of those numbers.

Found 3 solutions by checkley77, richwmiller, dabanfield:
Answer by checkley77(12844)

3x + 4y = 5 MULTIPLY BY 2
2x + 3y = 4 MULTIPLY BY -3 & ADD
6X+8Y=10
-6X-9Y=-12
--------------------
-Y=-2
Y=2 ans.
3X+4*2=5
3X+8=5
3X=5-8
3X=-3
X=-3/3
X=-1 ans.
Proof:
2*-1+3*2=4
-2+6=4
4=4

Answer by richwmiller(17219)

you aren't trying to eliminate numbers but letters.
for example
with the linear equations (they make a line on if you graph it)
x-y=7
2x+y=8
if you add the two equations y will be gone because we have a minus y and a plus y
with the equations
8x+y=23
3x+y=10
if we subtract one from the other we eliminate the y's because y-y =0
so you need to multiply your equations by a number such that you can subtract or add then and get rid of one of the letters
try multiplying the first one by 2 and the second one by 3
that way you'll have 6x in both and can subtract them.
then you solve for y
and then when you have y can plug y into any of the equations and solve for x
Me being a wise guy :-)
Did you miss class the day that was explained?







Answer by dabanfield(803)

3x + 4y = 5
2x + 3y = 4
Solve this problem using elimination.
If we multiply both sides of the first equation by 2 and both sides of the second equation by 3 we have:
6x + 8y = 10
6x + 9y = 12
We did this so we would have the same coefficient on x (that is 6). So now we can subtract the second equation from the first eliminating x:
(6-6)x +(8-9)y = 10-12
(0)x + (-1)y = -2
-y = -2
y = 2
Now substitute y=2 in either of the original equations and solve for x.